# Socialized Learning: A Survey of the Paradigm Shift for Edge Intelligence in Networked Systems

### v1

@@ IV-A1 The Main Role in the SL Architecture
In the SL, three key layers operate collaboratively. The Device Layer includes diverse end devices like drones, AR tools, health sensors, IoT devices, and smart home gadgets [54, 55, 56, 57]. These devices at the Device Layer, crucial for data production and collection, process data locally or offload it based on Quality of Experience (QoE) and Quality of Service (QoS) requirements. Edge Layer provides computing and storage, mainly through edge servers and gateways, orchestrating services and tasks for efficient data handling and resource optimization [58, 59]. Cloud Layer enhances traditional cloud computing with advanced capabilities for handling complex computations and large-scale storage, integrating closely with edge layer [60].
@@ IV-A2 Comparative Analysis in Different Aspects
Resource Allocation: Resource distribution differs notably among the three layers. The device layer, limited by its heterogeneity, has scarce resources, necessitating optimal algorithms for managing computation, storage, and energy [61]. While the edge layer has more resources, it’s still less capable than the cloud, emphasizing low-latency service and local data processing. The cloud layer, resource-rich, is mainly used for tasks demanding substantial computing strength.
@@
Task Execution: Task handling differs among the layers. The device layer manages simple, immediate tasks like data collection and preliminary analysis. The edge layer takes on tasks offloaded from devices, focusing on real-time actions with critical latency and energy concerns [63]. The cloud layer addresses complex tasks, such as in-depth data analysis and ML, leveraging its ample resources without tight time limits.
Computing Power: The device, edge, and cloud layers have varied computing strengths. The device layer is geared towards simpler tasks due to its limited power [64]. The edge layer, more powerful than devices, ensures swift processing of offloaded tasks but is still limited compared to the cloud. The cloud layer has abundant resources for intricate tasks but may lag in real-time processing because of its location. Each layer caters to distinct needs across applications.
@@ IV-B1 Non-Layer Architecture
IV-B1a Centralized Architecture
  • Resources are homogeneous across entities.
  • The processing capabilities of the entity are similar.
IV-B2 Intra-layer Architecture
@@ IV-B3 Inter-layer Architecture
The cloud-edge-device framework combines the computational strength of cloud computing with the edge’s low latency. This cloud-edge-device framework offers refined task and data distribution, thereby improving scalability and efficiency. While it can alleviate network bandwidth stress and reduce transmission latency, the edge’s limited computing capabilities for large-scale models can hinder its full potential. Research, such as the studies conducted by Morshed et al. analyzes the complexities of developing cloud-edge-device collaborative deep learning algorithms [80]. These algorithms are designed to be resource and data-aware, considering a multitude of factors like heterogeneity data models and resource availability. Yet, with an increase in the number of involved entities, focus optimization shifts to overall system cost and stability.
@@
Ad-hoc architecture, which allows for spontaneous resource distribution, is crucial in volatile or dynamic settings. It allows spontaneous resource distribution based on current system needs and limitations. In emergencies, resources might be allocated based on the event’s severity and available assets. Wang et al. suggest this strategy can be vital when traditional structures fall short due to fluctuating needs [81]. Ad-hoc architecture in SL offers flexibility for dynamic settings, but faces challenges in efficient resource allocation, balancing responsiveness with system efficiency. Despite these challenges, their adaptive nature renders them particularly relevant in certain contexts where responsiveness is paramount, like the unsupervised learning scenarios emphasized by Wang et al. [82].
Table V outlines attributes of architectures. Centralized architectures offer control and consistency, while intra- and inter-layer ones offer flexibility for dynamic needs. The appropriate architecture should be selected based on EI requirements.
@@ IV-C Lessons Learned
$w_{G},w_{i}$ Global and local model weights in federated learning
@@ V-A1 Mutual Assistance
Decentralized learning offers computing efficiency, scalability, and data locality, with SI building on these advantages to enhance privacy and resilience by maintaining data on local devices. These methods are notably robust against challenges such as data poisoning and system bottlenecks, depending on effective decentralized averaging and network topologies [23, 87]. Popular techniques for data aggregation in such environments include consensus-based algorithms and diffusion strategies [88].
@@
Figure 7: Illustration of DRL in Socialized Training.
@@ V-B1 Unidirectional Cooperation - Assistance
Split learning [93] exemplifies unidirectional assistance in inter-layer cooperation. Fig. 8 illustrates how the ML model is partitioned across multiple computational entities or ‘layers’. Devices process initial computation layers, forwarding outputs to edge servers for further processing. Devices benefit from edge servers’ computational capabilities. Shared features let edge servers process complex tasks while preserving data privacy and reducing communication overhead. Unlike approaches like FL, which update a global model across devices, it requires less data exchange.
@@
Figure 8: Illustration of Model Split Learning in Socialized Training.
Split learning suits scenarios requiring data privacy, like medical diagnosis [24], and in systems that require efficient use of computational resources. However, a challenge is managing the communication overhead from feature exchange in multi-device settings. The devices unilaterally assist the edge servers by undertaking a part of the overall computation, thereby allowing the assisted layer to perform more complex tasks or to do so more efficiently.
@@
In this context, the pioneering work by Lu et al. [95] introduce COLLA, a collaborative learning framework designed for user location prediction. In COLLA, the cloud-based model acts as a global aggregator, distilling insights from multiple edge-based models to create personalized models for each edge. The cloud-based model effectively assists edges in generating more accurate predictive models, embodying the spirit of unidirectional cooperation. Further extending this paradigm, Yao et al. [96] introduce a novel edge-cloud collaboration framework for recommendation systems that incorporates backbone-patch decomposition. This approach significantly reduces the computational burden on edges while maintaining high levels of personalization. The central model recalibrates edge backbone models to avoid local optimums through the MetaPatch mechanism.
@@ V-B2 Bidirectional Cooperation - Cooperation
where $BC$ is the collaborative outcome, $E$ and $C$ are the states of edge and cloud models, $D$ is the real-time edge data, $R$ is the cloud’s response, and $\theta$ are the collaboration parameters.
@@
Communication-Efficient Edge-Cloud Collaboration: Edge-cloud collaboration is a promising paradigm for leveraging the complementary advantages of edge and cloud computing to enable intelligent and responsive services. However, the communication between edge and cloud nodes can introduce significant overhead and latency, hindering the real-time performance of collaborative learning systems [104].
Researchers have explored communication-efficient strategies for edge-cloud collaboration, such as hierarchical compression and selective update transmission. These techniques aim to reduce the volume of data exchanged between edge and cloud nodes while preserving the quality of the learned models and the timeliness of the collaborative inference process. By optimizing the communication efficiency of edge-cloud collaboration, SL systems can harness the power of distributed computing resources while meeting the stringent latency and bandwidth constraints of real-world applications.
@@ V-B3 Bidirectional Cooperation - Exchange
One noteworthy application that captures this essence of exchange is the multi-modal federated collaborative training framework by Qayyum et al. [106]. The framework builds a composite model for diagnosing COVID-19 by harmoniously integrating two modalities of data-X-ray and ultrasound images - across different centers in a cloud-edge environment. Here, the interaction is not merely cooperative but fundamentally reciprocal, as each layer significantly contributes to the shared model and benefits from it for its own goals.
@@ V-C1 Conformity
Such a mechanism, by selectively amplifying certain portions of the input while downplaying others, embodies the principle of conforming to pivotal information [116]. Architectures like the Transformer, employing self-attention, underscore this conformity by weighting the importance of each word in relation to others. This evolution in design philosophy has ushered in models such as BERT [117] and GPT [118], which further epitomize ‘conformity’ by foregrounding contextually salient words in expansive text corpora. This mechanism is illustrated in Fig. 14. The SL function Eq. (1) for the attention mechanism is $F([\text{servers}],[\text{conformity}],\varnothing,\text{train},\text{data})$. The attention mechanism is:
@@ V-D1 Unidirectional Submissive Interaction - Suggestion&imitation
The diversity in data distribution across edges and the cloud accentuates the significance of heterogeneous transfer learning, as demonstrated in [123]. This learning approach adapts to varying data characteristics, employing symmetric transformations for domain-agnostic representations, as discussed in [124], and asymmetric transformations for domain-specific alignment, highlighted in [125]. This mirrors the social dynamic where one domain or group modulates its behavior to resemble another, often seen as superior. This concept is further exemplified by semi-supervised kernel matching domain adaptation [126], showcasing models’ capacity to assimilate and integrate more advanced insights.
@@
Submissive interaction in ML prominently manifests in three KD paradigms, each drawing parallels to social interactions. Response-Based KD imitates the direct responses of expert models, using these soft targets to guide student models. Feature-Based KD goes further, teaching student models to mimic intermediate features of teacher models, with FitNet as a prime example [129]. It aligns student and teacher feature maps for deeper knowledge transfer. Lastly, Relation-Based KD explores the relationships between layers, akin to the interplay of skills and concepts in social settings. The FSP matrix technique, capturing feature map pair relationships, is a notable method in this context [130]. Together, these KD methods reflect the ML journey from novice to expert, mirroring human learning processes.
@@
Meta learning, which aims to improve learning processes by utilizing experiences from past tasks, differs from transfer learning’s knowledge application, as shown in Fig. 17. Advances such as MELO [132], which extends MAML and applications involving spiking neural networks, demonstrate the adaptation of meta learning in EI. This evolution, when combined with model compression, equips models for the unique demands of EC [133].
@@ V-D2 Bidirectional Submissive Interaction
In machine societies, entities participate in bidirectional submissive interactions, meaning they both absorb and contribute knowledge. This dynamic allows them to switch roles between learner and mentor, ensuring that individual insights lead to collective system improvement and optimize overall efficiency.
Figure 17: Illustration of Meta Learning in Socialized Training.
@@
DA functions in a bidirectional or even multi-directional realm of transfer learning [134]. It leverages knowledge from a source domain to adapt to the target domain’s data distribution. This exchange is depicted in Fig. 18. Instance-based DA [135, 136] reweights source instances to bridge disparities between source and target samples. Feature-based DA [137, 138, 139] creates a unified space for source and target features to coalesce. For example, Gani et al.’s approach [139] utilizes GAN networks to create consistent features across domains.
@@
Insights from the target domain enhance the source’s knowledge. Liu et al. [140] describe this with their cycle self-training (CST) model, which facilitates knowledge exchange between domains. Such intricate dynamics of DA become indispensable in realms like EI where data distributions diversify across devices. Emblematic of this is the MobileDA framework by Yang et al. [141], which not only combats domain shifts in EI devices but also facilitates seamless knowledge exchange between centralized servers and edge entities.
@@
where $\mathcal{L}_{MTL}$ is the total multi-task loss, $\mathcal{L}_{t}$ is the loss for task $t$, $\alpha_{t}$ is the weight for task $t$, and $T$ is tasks’ total number.
Within the MTL, the categorization into hard sharing [143] and soft sharing [144] mirrors social dynamics where certain norms are universally rigid (hard sharing) while others permit regional variances (soft sharing). Maurer et al.’s exploration illustrates how shared features, envisioned as transformations through multi-layered networks [145], parallel the way societies adapt universal principles into specific regional customs. In the larger picture, MTL champions a symbiotic feature interaction across multiple tiers, harmonizing model structures and fostering task interdependencies. The upshot is enhanced model robustness, adaptability, and commendable prowess to tackle a plethora of complex tasks. Bidirectional submissive interactions, illustrated by DA and MTL, underscore the importance of knowledge exchange. For clarity, Table VIII provides a comparative analysis of these methods.
@@ V-E1 Competition
Competition is a driving force in many systems, pushing entities to strive for supremacy in various domains. In the machine society, algorithms and models mirror this competitiveness, aiming to outperform peers through constant fine-tuning, adaptation, and evolution. Each model aspires to set benchmarks for others to surpass.
@@
Ensemble learning epitomizes intra-layer competition where models vie to showcase their predictive prowess on a common dataset. While individual models are complete in themselves, their competitive edge is sharpened when juxtaposed with peers, pushing each model to accentuate its strengths and compensate for its weaknesses. This concept is illustrated in Fig. 20.
@@
Ensemble learning employs various strategies to harness this competitive spirit. In particular, approaches like voting [146] by pooling predictions, whereas weighted voting [147] confers preferential weights based on a model’s relative performance. Further, the model averaging method [148] seeks a consensus by averaging individual model outputs. Bai et al. [149] proposed an Automated DNN ensemble (AES), which selects optimal DNN models and tunes the ensemble according to nuances of the input data. Such intricate competitive dynamics have found success across many applications, from fraud detection [150] and medical diagnosis [151] to image recognition [152], emphasizing its effectiveness in harnessing intra-layer competition for superior predictive outcomes.

### v2

@@ IV-A1 The Main Role in the SL Architecture
In the SL, three key layers operate collaboratively. The Device Layer includes diverse end devices like UAVs, AR tools, health sensors, IoT devices, and smart home gadgets [57, 58, 59, 60]. These devices at the Device Layer, crucial for data production and collection, process data locally or offload it based on Quality of Experience (QoE) and QoS requirements. Edge Layer provides computing and storage, mainly through ESs and gateways, orchestrating services and tasks for efficient data handling and resource optimization [61, 62]. Cloud Layer enhances traditional cloud computing with advanced capabilities for handling complex computations and large-scale storage, integrating closely with edge layer [63].
@@ IV-A2 Comparative Analysis in Different Aspects
Resource Allocation: resource allocation differs notably among the three layers. The device layer, limited by its heterogeneity, has scarce resources, necessitating optimal algorithms for managing computation, storage, and energy [64]. While the edge layer has more resources, it’s still less capable than the cloud, emphasizing low-latency service and local data processing. The cloud layer, resource-rich, is mainly used for tasks demanding substantial computing power.
@@
Task Execution: Task handling differs among the layers. The device layer manages simple, immediate tasks like data collection and preliminary analysis. The edge layer takes on tasks offloaded from devices, focusing on real-time actions with critical latency and energy concerns [66]. The cloud layer addresses complex tasks, such as in-depth data analysis and ML, leveraging sufficient resources without tight time limits.
Computing Power: The device, edge, and cloud layers have varied computing capabilities. The device layer is geared towards simpler tasks due to its limited power [67]. The edge layer, more powerful than devices, ensures swift processing of offloaded tasks but is still limited compared to the cloud. The cloud layer has abundant resources for intricate tasks but may lag in real-time processing because of its location. Each layer caters to distinct needs across applications.
@@ IV-B1 Non-Layer Architecture
IV-B1a Centralized Architecture
  • Resources are homogeneous across entities.
  • The processing capabilities of the entity are similar.
IV-B2 Intra-layer Architecture
@@ IV-B3 Inter-layer Architecture
The cloud-edge-device framework combines the computing power of cloud computing with the edge’s low latency. This cloud-edge-device framework offers refined task and data distribution, thereby improving scalability and efficiency. While it can alleviate network bandwidth stress and reduce transmission latency, the edge’s limited computing capabilities for large-scale models can hinder its full potential. Research, such as the studies conducted by Morshed et al. analyze the complexities of developing cloud-edge-device collaborative DL algorithms [83]. These algorithms are designed to be resource and data-aware, considering a multitude of factors like heterogeneity data models and resource availability. Yet, with an increase in the number of involved entities, focus optimization shifts to overall system cost and stability.
@@
Ad-hoc architecture, which allows for spontaneous resource allocation, is crucial in volatile or dynamic settings. It allows spontaneous resource allocation based on current system needs and limitations. In emergencies, resources might be allocated based on the event’s severity and available assets. Wang et al. suggest this strategy can be vital when traditional structures fall short due to fluctuating needs [84]. Ad-hoc architecture in SL offers flexibility for dynamic settings but faces challenges in efficient resource allocation, balancing responsiveness with system efficiency. Despite existing the above challenges, their adaptive nature renders them particularly relevant in certain contexts where responsiveness is paramount, like the unsupervised learning scenarios emphasized by Wang et al. [85].
Table V outlines the attributes of architectures. Centralized architectures offer control and consistency, while intra- and inter-layer ones offer flexibility for dynamic needs. The selection of an appropriate architecture should be based on the specific requirements of the EI system.
@@ IV-C Lessons Learned
$w_{G},w_{i}$ Global and local model weights in FL
@@ V-A1 Mutual Assistance
Decentralized learning offers computing efficiency, scalability, and data locality, with SL building on these advantages to enhance privacy and resilience by maintaining data on local devices. These methods are notably robust against challenges such as data poisoning and system bottlenecks, depending on effective decentralized averaging and network topologies [23, 90]. Popular techniques for data aggregation in such environments include consensus-based algorithms and diffusion strategies [91].
@@
Figure 7: Illustration of distributed reinforcement learning in socialized training.
@@ V-B1 Unidirectional Cooperation - Assistance
Split learning [96] exemplifies unidirectional assistance in inter-layer cooperation. Fig. 8 illustrates how the ML model is partitioned across multiple computational entities or ‘layers’. Devices process initial computation layers, forwarding outputs to ESs for further processing. Devices benefit from ESs’ computational capabilities. Shared features let ESs process complex tasks while preserving data privacy and reducing communication overhead. Unlike approaches like FL, which update a global model across devices, it requires less data exchange.
@@
Figure 8: Illustration of split learning in socialized training.
Split learning is suitable for scenarios requiring data privacy, like medical diagnosis [24], and in systems that require efficient use of computational resources. However, a challenge is managing the communication overhead from feature exchange in multi-device settings. The devices unilaterally assist the ESs by undertaking a part of the overall computation, thereby allowing the assisted layer to perform more complex tasks or to do so more efficiently.
@@
In this context, the pioneering work by Lu et al. [98] introduces COLLA, a collaborative learning framework designed for user location prediction. In COLLA, the cloud-based model acts as a global aggregator, distilling insights from multiple edge-based models to create personalized models for each edge. The cloud-based model effectively assists edges in generating more accurate predictive models, embodying the spirit of unidirectional cooperation. Further extending this paradigm, Yao et al. [99] introduce a novel edge-cloud collaboration framework for recommendation systems that incorporates backbone-patch decomposition. This approach significantly reduces the computational burden on edges while maintaining high levels of personalization. The central model recalibrates edge backbone models to avoid local optimums through the MetaPatch mechanism.
@@ V-B2 Bidirectional Cooperation - Cooperation
where $BC$ is the collaborative outcome, $E$ and $C$ are the states of edge and cloud models, $D$ is the real-time edge data, $R$ is the cloud’s response, and $\theta$ is the collaboration parameters.
@@
Communication-efficient Edge-cloud Collaboration: Edge-cloud collaboration is a promising paradigm for leveraging complementary advantages of edge and cloud computing to enable intelligent and responsive services. However, the communication between edges and the cloud can introduce significant overhead and latency, hindering real-time performance of collaborative learning systems [107].
Researchers have explored communication-efficient strategies for edge-cloud collaboration, such as hierarchical compression and selective update transmission. These techniques aim to reduce the volume of data exchanged between edge and cloud nodes while preserving the quality of the learned models and the timeliness of the collaborative inference process. Optimizing edge-cloud collaboration’s communication efficiency allows SL systems to harness distributed computing resources while meeting real-world applications’ stringent latency and bandwidth constraints.
@@ V-B3 Bidirectional Cooperation - Exchange
One noteworthy application that captures this essence of exchange is the multi-modal federated collaborative training framework by Qayyum et al. [109]. The framework builds a composite model for diagnosing COVID-19 by harmoniously integrating two modalities of data-X-ray and ultrasound images, across different centers in a cloud-edge environment. Here, the interaction is not merely cooperative but fundamentally reciprocal, as each layer significantly contributes to the shared model and benefits from it for its own objectives.
@@ V-C1 Conformity
Such a mechanism, by selectively amplifying certain portions of the input while downplaying others, embodies the principle of conforming to pivotal information [119]. Architectures like the Transformer, employing self-attention, underscore this conformity by weighting the importance of each word in relation to others. This evolution in design philosophy has ushered in models such as BERT [120] and GPT [121], which further epitomize ‘conformity’ by foregrounding contextually salient words in expansive text corpora. This mechanism is illustrated in Fig. 14. The SL function Eq. (1) for the attention mechanism is $F([\text{servers}],[\text{conformity}],\varnothing,\text{train},\text{data})$. The attention mechanism is defined as follows:
@@ V-D1 Unidirectional Submissive Interaction - Suggestion&imitation
The diversity in data distribution across edges and the cloud highlights the significance of heterogeneous transfer learning, as demonstrated in [126]. This learning approach adapts to varying data characteristics, employing symmetric transformations for domain-agnostic representations, as discussed in [127], and asymmetric transformations for domain-specific alignment, highlighted in [128]. This mirrors the social dynamic where one domain or group modulates its behavior to resemble another, often seen as superior. This concept is further exemplified by semi-supervised kernel matching domain adaptation [129], showcasing models’ capacity to assimilate and integrate more advanced insights.
@@
Submissive interaction in ML prominently manifests in three KD paradigms, each drawing parallels to social interactions. Response-based KD imitates the direct responses of expert models, using these soft targets to guide student models. Feature-based KD goes further, teaching student models to mimic intermediate features of teacher models, with FitNet as a prime example [132]. It aligns student and teacher feature maps for deeper knowledge transfer. Lastly, Relation-based KD explores the relationships between layers, akin to the interplay of skills and concepts in social settings. The feature space projection matrix, capturing feature map pair relationships, is a notable method in this context [133]. Together, these KD methods reflect the ML trends from novice to expert, mirroring human learning processes.
@@
Meta learning, which aims to improve learning processes by utilizing experiences from past tasks, differs from transfer learning’s knowledge application, as shown in Fig. 17. Advances such as MELO [135], which extends model-agnostic meta-Learning (MAML) and applications involving spiking neural networks, demonstrate the adaptation of meta learning in EI. This evolution, when combined with model compression, equips models for the unique demands of EC [136].
Figure 17: Illustration of meta learning in socialized training.
@@ V-D2 Bidirectional Submissive Interaction
In machine societies, entities participate in bidirectional submissive interactions, meaning they both acquire and contribute knowledge. This dynamic allows them to switch roles between learner and mentor, ensuring that individual insights lead to collective system improvement and optimize overall efficiency.
@@
DA functions in a bidirectional or even multi-directional realm of transfer learning [137]. It leverages knowledge from a source domain to adapt to the target domain’s data distribution. This exchange is depicted in Fig. 18. Instance-based DA [138, 139] reweights source instances to bridge disparities between source and target samples. Feature-based DA [140, 141, 142] creates a unified space for the cohesion of source and target features. Gani et al.’s approach [142] utilizes generative adversarial networks (GANs) to create consistent features across domains.
@@
Insights from the target domain enhance the source’s knowledge. Liu et al. [143] describe this with their cycle self-training (CST) model, which facilitates knowledge exchange between domains. Such intricate dynamics of DA become indispensable in realms like EI where data distributions diversify across devices. Representative of this is the MobileDA framework by Yang et al. [144], which not only combats domain shifts in EI devices but also facilitates seamless knowledge exchange between centralized servers and edge entities.
@@
where $\mathcal{L}_{MTL}$ is the total multi-task loss, $\mathcal{L}_{t}$ is the loss for task $t$, $\alpha_{t}$ is weight for task $t$, and $T$ is total number of tasks.
Within the MTL, the categorization into hard sharing [146] and soft sharing [147] mirrors social dynamics where certain norms are universally rigid (hard sharing) while others permit regional variances (soft sharing). Maurer et al.’s exploration illustrates how shared features, envisioned as transformations through multi-layered networks [148], parallel the way societies adapt universal principles into specific regional customs. In the larger picture, MTL champions a symbiotic feature interaction across multiple tiers, harmonizing model structures and fostering task interdependencies. The result is enhanced model robustness, adaptability, and commendable prowess to tackle a plethora of complex tasks. Bidirectional submissive interactions, illustrated by DA and MTL, underscore the importance of knowledge exchange. For clarity, Table VIII provides a comparative analysis of these methods.
@@ V-E1 Competition
Competition is a driving force in many systems, pushing entities to strive for supremacy in various domains. In the machine society, algorithms and models mirror this competitiveness, aiming to outperform peers through constant fine-tuning, adaptation, and evolution. Each model desires to set benchmarks for others to surpass.
@@
Ensemble learning epitomizes intra-layer competition where models vie to showcase their predictive prowess on a common dataset. While individual models are complete in themselves, their competitive advantages become more pronounced when juxtaposed with peers, pushing each model to highlight its strengths and compensate for its weaknesses. This concept is illustrated in Fig. 20.
@@
Ensemble learning employs various strategies to harness this competition. In particular, approaches like voting [149] by pooling predictions, whereas weighted voting [150] confers preferential weights based on a model’s relative performance. Further, the model averaging method [151] seeks a consensus by averaging individual model outputs. Bai et al. [152] proposed an Automated DNN ensemble (AES), which selects optimal DNN models and tunes the ensemble according to nuances of the input data. Such intricate competitive dynamics have found success across many applications, from fraud detection [153] and medical diagnosis [154] to image recognition [155], emphasizing its effectiveness in harnessing intra-layer competition for superior predictive outcomes.